For all problems assume an effective monthly interest rate of 1% unless otherwise indicated in the problem.

3. In the CME’s silver futures market, the current silver price is $33.41 per troy ounce, and the contract size is 5000 troy ounces.

(a) If a hedge fund wanted to be long $835,250 worth of silver futures, how many contracts should they buy?

(b) If the CME requires an 12% initial margin, how much does the hedge fund need to post in the margin account?

(c) If the price of silver closes tomorrow at $33.50 per troy ounce, what is the balance in the margin account?
Answer #1

Answer (a) Total Contract Value = USD 835,250

Contract Size = 5,000 Troy Ounces

Contract Value = 5,000 * 33.41 = USD 167,050

No. of Contract required to be purchased = USD 835,250 / USD 167,050 = 5 Contracts

Answer (b) Initial Margin = 12% of USD 835,250 = USD 100,230

Answer (c) Total Value of Contract = USD 835,250

Less: Margin Money = (USD 100,230)

Balance = USD 735,020

Interest on Balance at 1% for 1 day = USD 735,020 * 1% / 30 days = USD 245

Computation of Margin Balance:

Initial Margin = USD 100,230

Add: Profit on contract = USD 2,250

(33.5 - 33.41) * 5000 * 5

Less: Interest chargable = (USD 245)

Closing Balance on Margin account = USD 102,235